### v2

@@ 4.2 Preliminary simulations
All six RL approaches were initially trained against five agents who made random bids at or below their item values. A total of 54 training sessions were performed with the resultant model (or Q-table) saved. For the six algorithms, training runs were performed using 100,000 episodes for each combination of auction type (DP, GSP and UP) and number of items on offer (4, 6 and 8).
@@ 4.3 Simulation results
We begin with the learning results for the three auctions in three different scenarios. Figure 1 shows the learning ratios of the six algorithms for the Discriminatory Price auction across three scenarios. As shown in this figure, Proximal Policy Optimization (PPO) is the most stable algorithm in terms of the learning ratio and is, in some sense, the fastest. However, all six algorithms demonstrate reasonable convergence to similar strategies by the final episodes. It should be noted that since bids are ranked from highest to lowest, the order in which an agent submits their highest bids (first or second) does not impact the outcome. In all cases, agents bid below their values, a straightforward conclusion from the theory (Krishna, 2009). Moreover, when more units are available, agents tend to reduce their second bids further, which is also very intuitive.
@@
A similar pattern of learning is observed in Figure 2 for the Generalized Second Price auction. Again, in all cases, the agents bid below their values, aligning with theoretical predictions (Edelman et al., 2007). The DPN algorithm struggles to learn an optimal strategy, and as the number of items increases, its performance worsens. This is mainly because DPN is significantly influenced by stochastic outcomes, and in the GSP format, where the payoff is determined by someone else’s bid, DPN struggles to find a consistent learning path. This issue is most pronounced in the Uniform Price auction, where computing an expectation for the highest losing bid from a probabilistic perspective is much more challenging.
Finally, Figure 3 illustrates the learning patterns for the Uniform Price auction. In PPO, Actor-Critic (A2C), and Q-learning, when the number of items is low, bids are less truthful. While Q-learning maintains this strategy even as the number of items increases, PPO and A2C shift towards more truthful bids with more items available. We later show that this adjustment allowed these two algorithms to achieve the highest payoffs among all agents.
Payoffs
@@
Having shown the learning patterns of all agents, we next demonstrate their performance in terms of payoffs, which is the most important metric for their behavior. As illustrated in Tables 2, 3, and 4, PPO consistently achieves the highest payoff among the six algorithms without any exceptions. A2C is the second-best performing algorithm in almost all cases, except for two instances where VPG ranks second. The other algorithms vary in terms of their payoff rankings across the three different scenarios and auctions. Notably, Q-learning, despite its simplistic structure, ranks among the top three algorithms in several scenarios, including DP and GSP with four units and DP with eight units.
Auction performance
@@
Next, we show how each auction performed in terms of revenue and efficiency. Figure 4 depicts the revenue and efficiency for each auction across the three scenarios. Focusing on later episodes, where the learning patterns stabilize, we observe that with four items available, all three auctions generate similar revenues, with the Uniform Price (UP) auction slightly outperforming the others. However, as the number of units increases, this ranking changes; with eight items available, the Discriminatory Price (DP) auction generates the highest revenue, followed by the Generalized Second Price (GSP) and then UP. In terms of efficiency, UP dominates the other two auctions in all scenarios, followed by GSP and DP.
@@ 4.3.1 Simulation results with all PPO agents
Given the results above, it is clear that the PPO agent significantly outperforms the others in terms of learning to bid optimally. However, observations of the learning patterns indicate that the auction ranking in terms of revenue and efficiency is influenced by other sub-optimal learning behaviors. It is worthwhile to explore how each auction performs when all six agents are PPO. Figure 5 shows the results regarding the revenue and efficiency rankings of the three auctions when all agents are PPO. The Uniform-Price auction remains the most efficient among the three, closely followed by GSP. However, the revenue rankings are reversed, with GSP now generating the highest revenue, followed by UP.
Two clear conclusions can be drawn from these observations: first, the Uniform-Price auction consistently performs well in terms of efficiency; second, GSP is the most stable auction in terms of both revenue and efficiency, regardless of whether the agents bid optimally or not.

### v3

@@ 4.2 Preliminary simulations
The six RL approaches were initially trained against five agents who made random bids at or below their item values. A total of 54 training sessions were performed with the resultant model (or Q-table) saved. For the six algorithms, training runs were performed using 100,000 episodes for each combination of auction type (DP, GSP and UP) and number of items on offer (4, 6 and 8).
@@ 4.3 Simulation results
In this subsection, we report the simulation results. We begin by examining the learning ratio for each auction and discuss variations in learning and bidding strategies as the number of available items changes. We then analyze the payoffs that various algorithms achieved in each auction, using these as measures of each algorithm’s bidding performance. Subsequently, we focus on the revenue and efficiency of each auction to assess overall performance.
4.3.1 Learning
@@
We begin with the learning results for the three auctions in three different scenarios with respect to the number of items available. Figure 1 shows the learning ratios of the six algorithms for the discriminatory price auction in three scenarios. As shown in this figure, PPO is the most stable algorithm in terms of the learning ratio and is, in some sense, the fastest. However, all six algorithms demonstrate reasonable convergence to similar strategies by the final episodes. It is important to note that since bids are ranked from highest to lowest, the order in which an agent submits their highest bids (first or second) does not impact the outcome. In all cases, agents submit bids below their values, a straightforward conclusion from the theory, that is, they do not play a dominated strategy (Krishna, 2009). Moreover, when more units are available, agents tend to reduce their second bids further, which is also very intuitive.
@@
A similar pattern of learning is observed in Figure 2 for the GSP auction. In all cases, agents bid below their values, which supports that truthful bidding is not optimal in GSP (Edelman et al., 2007). One interesting observation is that the PPO agent submits bids that are very similar to the ones submitted by the same agent in the DP auction. We later show that these bidding strategies resulted in PPO becoming the best performing algorithm in both auctions. The DPN algorithm struggles to learn an optimal strategy and as the number of items increases, its performance worsens. This is mainly because DPN is significantly influenced by stochastic outcomes, and in the GSP format, where the payoff is determined by someone else’s bid, DPN struggles to find a consistent learning path. This issue is most pronounced in the Uniform Price auction, where computing an expectation for the highest losing bid from a probabilistic perspective is much more challenging.
Finally, Figure 3 illustrates the learning patterns for the uniform price auction. When the number of items is at its lowest, PPO and A2C submit bids that are less truthful, with the level of truthful bidding increasing as the number of items increases to six. However, we observe overbidding with both PPO and A2C when there are eight items available. While this may seem unintuitive, as bidding equal to the value is a weakly dominant strategy in the UP auction (Krishna, 2009), the likelihood of incurring a loss from overbidding declines as the number of items increases and could potentially become very low. Thus, the strategies that PPO and A2C agents follow would not necessarily result in lower payoffs compared to other agents. In fact, we later show that the strategies that PPO and A2C followed in UP allowed these two algorithms to achieve the highest payoffs among all agents. Moreover, Q-learning maintains an underbidding strategy across all cases, even increasing the extent of underbidding as the number of items increases while VPG starts with almost truthful bids for both units and underbids more when the total number of items increase. Overall, we observed the highest level of variation in strategies for the UP auction, which shows that despite its simple structure, optimal bidding strategies in this auction are not easy to learn.
4.3.2 Payoffs
@@
Having shown the learning patterns of all agents, we next demonstrate their performance in terms of payoffs, which is the most important metric for their behavior. As illustrated in Tables 2, 3, and 4, PPO consistently achieves the highest payoff among the six algorithms without any exceptions. A2C is the second-best performing algorithm in almost all cases, except for two instances where VPG ranks second. The other algorithms vary in terms of their payoff rankings across the three different scenarios and auctions. In particular, Q-learning, despite its simplistic structure, is among the top three algorithms in several scenarios, including DP and GSP with four units and DP with eight units.
4.3.3 Auction performance
@@
Next, we show how each auction performed in terms of revenue and efficiency. Figure 4 depicts the revenue and efficiency for each auction across the three scenarios. Focusing on later episodes, where the learning patterns stabilize, we observe that with four items available, all three auctions generate similar revenues, with the UP auction slightly outperforming the others. However, as the number of units increases, this ranking changes; with eight items available, the DP auction generates the highest revenue, followed by the GSP and then UP. In terms of efficiency, UP dominates the other two auctions in all scenarios, followed by GSP and DP.
@@ 4.3.4 Simulation results with all PPO agents
Given the results above, it is clear that the PPO agent significantly outperforms the others in terms of learning to bid optimally. However, observations of the learning patterns indicate that the auction ranking in terms of revenue and efficiency is influenced by other sub-optimal learning behaviors. It is worthwhile to explore how each auction performs when all six agents are PPO. Figure 5 shows the results regarding the revenue and efficiency rankings of the three auctions when all agents are PPO. The UP auction remains the most efficient among the three, closely followed by GSP. However, the revenue rankings are reversed, with GSP now generating the highest revenue, followed by UP.
Two clear conclusions can be drawn from these observations: first, the uniform price auction consistently performs well in terms of efficiency; second, GSP is the most stable auction in terms of both revenue and efficiency, regardless of whether the agents bid optimally or not.